A semicircular window of radius 16 inches is to be laminated with a sunblock coating that costs $0.65 per square inch to apply. What is the total cost of coating the window, to the nearest cent? (Use п = 3.14. Hint: Area of Circle: A = пr2)
Answer by checkley77(12844)

A=PIR^2/2
A=3.14*16^2/2
A=3.14*256/2
A=3.14*128
A=401.92 IN^2
401.92*.65=@261.25 ANS.